Session: Feb 2004 (session 1)
Wavelength: 1.3 cm
To give feedback on this experiment, click
Medicina (Monari at Wed Feb 11 21:59:24 2004) Success :
No problem durinf the observation
Effelsberg (A. Kraus at Wed Feb 11 23:01:22 2004) Success :
Noto (P.Cassaro at Thu Feb 12 00:45:00 2004) Success :
Shanghai (Xinyong Huang at Thu Feb 12 03:29:20 2004) Success :
Metsahovi (Ari Mujunen at Thu Feb 12 07:21:43 2004) Minor failures : PCAL
No pcal connected.
Onsala (Michael Lindqvist at Thu Feb 12 08:34:07 2004) Minor failures :
Problems to get any level on VC08. Could have been a bad cable.
Jodrell_Bank (Alastair Gunn at Mon Feb 16 12:40:53 2004) Success :
On disk. Repatched only BBCs 5,6,7,8 to high. No known problems.
Cambridge (Alastair Gunn at Mon Feb 16 12:41:30 2004) Success :
On tape. Remapped system tracks to avoid bad track. Signal only
in BBC01 due to MERLIN link restriction. No known problems.
Urumqi (Liu Xiang at Fri Feb 20 12:50:27 2004) Minor failures :
The LO had shifted to 22000MHz we checked after the exp.